OpenStudy (anonymous):

what is the difference of 9/10 - 1/2

OpenStudy (gunsodo):

4/10 = 2/5 is the answer

OpenStudy (gunsodo):

Change 1/2 to 5/10 by multiplying the dominators and numerators by 5 You'll get 9/10 - 5/10 = 4/10

OpenStudy (anonymous):

he is right :)